OpenStudy (anonymous):

Is it possible to verify: (sin theta/1 + cos theta) +(1+cos theta/ sin theta)= 2csc theta.

OpenStudy (kainui):

Yes. Try your best to solve it and show me where you get stuck. Also, I suggest using more parentheses so that it's easier for me to read. So maybe like this: [sin theta /(1+cos theta)] + ...

OpenStudy (anonymous):

\[[(\sin \theta)/(1+\cos \theta)]* [(1-\cos \theta/\sin \theta)] +[(1+\cos \theta)/(\sin \theta)]*[(\sin \theta)/(1-\cos \theta)]\] ?

zepdrix (zepdrix):

\[\Large\rm \frac{\sin \theta}{1+\cos \theta}+\frac{1+\cos \theta}{\sin \theta}\]Hmmmm not sure what you were trying to do there >.< If you're going to multiply top and bottom by something, they need to be the same. Example: (1-cosx)/(1-cosx). Try to remember conjugates! You're usually trying to get back to your Pythagorean Identity whenever you see 1+cos or 1+sin

zepdrix (zepdrix):

\[\Large\rm \frac{\sin \theta}{1+\cos \theta}\color{royalblue}{\left(\frac{1-\cos \theta}{1-\cos \theta}\right)}+\frac{1+\cos \theta}{\sin \theta}\]I would probably give the first fraction this, the conjugate of the bottom. It should help clean things up.

OpenStudy (anonymous):

Oh okay, sorry. Thank you for clearing that up for me! I'll try reworking it now.